OpenStudy (anonymous):

Solve the system of equations by the substitution method. If the system does not have a single ordered pair as a solution, state whether the system is inconsistent or dependent. x + 3y = 32 -3x + 2y = 3 A. (4, 10) B. No solution; inconsistent C. (5, 9) D. (-5, 10)

jimthompson5910 (jim_thompson5910):

first solve x + 3y = 32 for x x + 3y = 32 x = 32 - 3y then plug this into -3x + 2y = 3 -3x + 2y = 3 -3(32 - 3y) + 2y = 3 ... plug in x = 32 - 3y -96 + 9y + 2y = 3 11y - 96 = 3 I'll let you finish up. You need to solve that for y. Once you've done that, you use the value of y to find x

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

c?

jimthompson5910 (jim_thompson5910):

correct, nice work